OpenStudy (anonymous):

Solve the inequality. 3m 9 Solve the inequality. 3m 9 @Mathematics

jimthompson5910 (jim_thompson5910):

something is missing....

OpenStudy (anonymous):

3m is greater than or equal to negative 9

jimthompson5910 (jim_thompson5910):

ok thx

jimthompson5910 (jim_thompson5910):

\[\Large 3m \ge -9\] \[\Large m \ge \frac{-9}{3}\] \[\Large m \ge -3\]